OpenStudy (jiteshmeghwal9):

Prove :- \[2+2=2\times2\]

hartnn (hartnn):

both sides =4

Parth (parthkohli):

Intuition, man, intuition.

Parth (parthkohli):

Okay, so do you know that...\[\rm a + a + a\cdots(b~times) = a\times b\]?

OpenStudy (jiteshmeghwal9):

ok ! @hartnn why \(2+2 = 2 \times 2=4\) ?

OpenStudy (jiteshmeghwal9):

yeah @ParthKohli :)

Parth (parthkohli):

As I told you,\[\rm 2 + 2 = 2 + 2\cdots(2~times) = 2 \times 2\]

Parth (parthkohli):

It's not a formal proof, but pretty decent.

OpenStudy (jiteshmeghwal9):

ok ! so why this case is not with other numbers such as why \(3+3\neq 3\times3\)

OpenStudy (anonymous):

because when you have 2x2, it means 2+2. if you have 3x3, it means 3+3+3, not 3+3. this also works with zero (0x0=0+0).

OpenStudy (jiteshmeghwal9):

but my question is why 2*2 means 2+2 @etemplin :)

OpenStudy (anonymous):

multiplication is a shorthand for of addition when you have some number n, n+n+n+n, can be rewritten as a product 4n so for 2x2, you can rewrite it as 2+2 which = 2+2 = 4

OpenStudy (jiteshmeghwal9):

hmm...

OpenStudy (anonymous):

it just happens to work that 2x2=2+2. theres nothing special about it

OpenStudy (anonymous):

multiplication just tells you how much is adding, like 3x3 means you add 3, 3 times and so for 4n, it means you add n 4 times so for 2x2 it means you add 2, 2 times

OpenStudy (jiteshmeghwal9):

nice reply all of u :) Thanx @jayz657 :)

OpenStudy (anonymous):

np glad i could help

OpenStudy (jiteshmeghwal9):

:)